OpenStudy (anonymous):

Michelle Payne deposited $8,000 in a savings account paying 6.25% simple interest. How long will it take for her investment to amount to $10,000?

OpenStudy (imstuck):

what formula are you using for this? The one I have is this:\[A=A _{0}(1+r)^{t}\]Is that the one you learned for this type of problem?

OpenStudy (imstuck):

\[10,000=8,000(1+.0625)^{t}\]and solve for t

OpenStudy (tkhunny):

Simple Interest. $10,000 - $8,000 = $2,000 \(2000.00 = 8000.00 \cdot 0.0625 \cdot t\)